how to calculate the theoretical moment of inertia of disk. mass is 68.9 grams and diameter is 10.3 cm. use equation I=1/2mr^2
Answer #1

I = mr^2 / 2

m = 68.9 g = 68.9 * 10^(-3) kg

r = 10.3/2 cm = 5.15 *10^(-2) m

I = 68.9*10^(-3 ) * (0.0515)^2 / 2 =0.0914 * 10^(-3) kg m^2 = 91.4 *10^(-6) kg m^2